# Versor: A Geometric Sequence Architecture Enhanced Scale Generalization and Interpretability via Conformal Algebra

### v1

@@ Abstract
A novel sequence architecture design is introduced, Versor, which uses Conformal Geometric Algebra (CGA) in place of the traditional fundamental non-linear operations to achieve structural generalization and significant performance improvements on a variety of tasks, while offering improved interpretability and efficiency. By embedding states in the $\mathcal{C}l_{4,1}$ manifold and evolving them via geometric transformations (rotors), Versor natively represents $SE(3)$-equivariant relationships without requiring explicit structural encoding.
Versor is validated on chaotic N-body dynamics, topological reasoning, and standard multimodal benchmarks (CIFAR-10, WikiText-103), consistently outperforming Transformers, Graph Networks, and geometric baselines (GATr, EGNN). Key results include: orders of magnitude fewer parameters ($200\times$ vs. Transformers); interpretable attention decomposing into proximity and orientational components; zero-shot scale generalization (99.3% MCC on topology vs. 50.4% for ViT); and $O(L)$ linear complexity via the novel Recursive Rotor Accumulator. In out-of-distribution tests, Versor maintains stable predictions while Transformers fail catastrophically. Custom Clifford kernels achieve up to $78\times$ speedup, providing a scalable foundation for geometrically-aware scientific modeling.
@@ 1 Introduction
The remarkable success of the Transformer architecture [28] has cemented the “Sequence of Vectors” paradigm as the standard for artificial intelligence. Whether the modality is text, images (ViT [10]), or audio, data is tokenized and projected into a flat, high-dimensional Euclidean space ($\mathbb{R}^{d_{model}}$). In this space, the relationships between features are modeled using the dot product $\mathbf{q}^{T}\mathbf{k}$, a scalar measure of similarity.
However, the physical world is not merely a collection of features, but a realization of physical laws acting on a structured manifold. Standard neural networks also treat data as points in a flat Euclidean space $\mathbb{R}^{d}$, relying on the dot product (a measure of similarity) as the primary relational primitive. This approach is geometrically naive: it ignores the fundamental symmetries of our universe (rotation, translation, and scaling). To respect these symmetries, a standard Transformer must expend vast computational resources to “learn” invariants from millions of augmented examples – a process that approximates what could be analytically enforced by a simple algebraic group action [2]. This fundamental mismatch between the geometry of the world and the architecture of the model is termed the “Euclidean Bottleneck”.
This work argues that the next leap in AI Structural Intelligence requires embedding these symmetries directly into the substrate of the network. Versor is presented as an architecture built not on linear algebra but on Conformal Geometric Algebra (CGA). Specifically, the architecture utilizes $\mathcal{C}l_{4,1}$, a 5-dimensional framework that linearizes the conformal group of 3D Euclidean space.
@@ 1.1 Key Contributions
    Scale Generalization: Demonstration that geometric priors enable scale-invariant reasoning—Versor achieves 99.3% MCC on topological connectivity tasks (vs. 50.4% for Vision Transformers) and maintains structural performance across varying sequence densities.
@@
    Hardware-Efficient Implementation: Custom Triton and MLX kernels using bit-masked basis contraction to compute Clifford products, achieving up to $78\times$ speedup and eliminating the memory bottleneck (typically $50\times$ reduction) of standard sparse implementations.
@@ 2 Background: Transformers, CGA, and the Manifold Hypothesis
Transformers and the Euclidean Bottleneck: Artificial Intelligence is dominated by the Transformer [28]. While effective, its reliance on flat Euclidean space ($\mathbb{R}^{d}$) forces it to learn symmetries ($SE(3)$) from data rather than enforcing them algebraically [2]. This is termed the “Euclidean Bottleneck.”
Conformal Geometric Algebra ($\mathcal{C}l_{4,1}$): The architecture is constructed on $\mathcal{C}l_{4,1}$, a 32-dimensional algebra generated by $\{e_{1},e_{2},e_{3},e_{+},e_{-}\}$. This framework isometrically lifts 3D points $\mathbf{x}$ to null vectors $X$ in 5D space:
@@
This ensures that distance calculations are linearized (see Appendix C.7). Crucially, transformations are represented uniformly as rotors $R$, which act on state vectors $\Psi$ via the sandwich product $\Psi^{\prime}=R\Psi\widetilde{R}$. This structure enforces the Manifold Hypothesis: by limiting latent states to the Spin group, $\text{Spin}(4,1)\subset\mathcal{C}l^{+}_{4,1}$, we guarantee valid physical transformations (isometries), preventing unphysical shearing.
For a detailed introductory treatment of Clifford algebra and their applications, we refer to [8], see Appendix A for explanations of the basis relations and multivector structure essential to this work.
@@ 3 Related Work
Geometric Deep Learning attempts to encode symmetry priors into neural networks [2], with recent focus on $SO(3)$-equivariant CNNs [6] and the Geometric Algebra Transformer (GATr) [1]. Unlike static-processing baselines like GATr and CGENN [22] which remain “frame-centric” and require $O(L^{2})$ attention, Versor focuses on path-based sequence modeling. By evolving a latent spinor via the Recursive Rotor Accumulator (RRA), Versor captures temporal coherence as a continuous trajectory on a Lie manifold, providing an inductive bias for dynamical systems that standard frame-based models lack.
This approach aligns with sub-quadratic architectures like Mamba [12] and RWKV [20], but provides a purely geometric interpretation of the hidden state. While concurrent methods like CARE [25] utilize Clifford rotors for positional encodings in Euclidean Transformers, Versor is a complete sequence architecture defined strictly within $\mathcal{C}l_{4,1}$, achieving $O(L)$ linear scaling. Furthermore, by embedding dynamics directly into the manifold, Versor avoids the need for the manually-enforced governing equations required by PINNs [21] and HNNs [11], while achieving better long-horizon stability than GNS [23].
@@ 4 The Versor Architecture
Versor is a sequence architecture that replaces the vector-space assumptions of Transformers with the graded manifold structure of $\mathcal{C}l_{4,1}$. The model processes a stream of multivectors via two core mechanisms: Geometric Product Attention (GPA) for relational reasoning and the Recursive Rotor Accumulator (RRA) for temporal integration.
@@ 4.2 Component 2: Recursive Rotor Accumulator (RRA)
Manifold Normalization: Crucially, the constrained $\Psi\widetilde{\Psi}=1$ is enforced at every step. This projects numerical drifts back into the manifolds, acting as a geometric regularizer that prevents the ”exploding state” problem of standard RNNs (see Appendix B for stability proofs.)
Hamiltonian Extension: For strict energy conservation, a Hamiltonian-Versor Hybrid is optionally deployed where the network parameterizes an energy surface $H(q,p)$ rather than predicting steps directly (details in Appendix F.1).
@@ 6 Experiments
The Versor architecture is now evaluated against state-of-the-art baselines across a variety of challenging tasks: chaotic N-body dynamics, topological reasoning, and multimodal learning (NLP, Vision, Graph). Across all benchmarks, Versor consistently outperforms comparative architectures in terms of accuracy, generalization, and parameter efficiency, validating the effectiveness of geometric inductive biases for sequence modeling.
@@ 6.1.1 Qualitative Geometric Intelligence
The Geometric Product Attention (GPA) is decomposed into scalar and bivector components (Figure 2). Scalar maps recover distance-based interaction laws, while bivector intensity highlights torque-maximizing interactions, validating that Versor attends to the full geometric configuration.
@@ 6.1.2 Long-Horizon Energy Stability
Models are rolled out for $T=50$ steps. The deviation of total system energy $H=T+V$ (kinetic plus potential) is measured as percentage drift. Conservation of energy is a fundamental physical constraint; a model that has truly learned the underlying dynamics should approximately conserve energy. In our parity tests, we observe that while large-scale models (e.g. 2.2M GATr) can achieve near-perfect conservation, their performance degrades significantly when matched to Versor’s minimal parameter count (6.6K), highlighting the strength of Versor’s geometric inductive bias.
Table 1: Performance on Static 5-Body Dynamics (Comparison with State-of-the-Art). Mean $\pm$ std over 5 random seeds. Versor achieves superior accuracy (5.21 MSE) compared to both Graph Networks and Transformers while maintaining extreme parameter efficiency.
@@
Transformer ($d=128$) 1.320M 0.82 $6.609\pm 6.415$ $381.1\pm 41.1$
Transformer (Large) 37.8M 24.88 $3.120\pm 1.250$ $210.5\pm 35.2$
Mamba 0.016M 0.05 $19.930\pm 7.850$ $3938\pm 2623$
GNS [23] 0.026M 0.31 $5.881\pm 6.408$ $366.7\pm 85.7$
HNN [11] 0.021M 0.12 $4.826\pm 6.378$ $10.7\pm 1.1$
EGNN [24] 0.030M 0.32 $6.695\pm 5.936$ $723.9\pm 351.2$
GATr 0.018M 1.29 $0.680\pm 7.070 11$$16.8\pm 56.1$
Versor (Base) 0.006M 1.20 $\mathbf{5.210\pm 6.387}$ $\mathbf{133.0\pm 37.7}$
Versor (Multi-Channel) 1.1M 5.45 $\mathbf{3.067\pm 0.14}$ $\mathbf{144.3\pm 0.0}$
Ham-Versor 0.044M 4.24 $4.827\pm 6.379$ $2.4\pm 1.7$
* Latency measured on CPU as per-step inference. ‡Model downscaled to match Versor’s parameter scale for inductive bias assessment. For fair architectural comparison, GATr was accelerated using our optimized multivector kernel.
The 4-channel variant results are detailed in Appendix F.1.4.
@@
Figure 3: Computational Scaling: Latency (ms) vs. Sequence Length $L$. Versor maintains strictly linear $O(L)$ growth, whereas standard Transformers diverge quadratically, reaching memory exhaustion (OOM) at $L=1024$. The dotted reference line explicitly denotes a gradient of 1 (linear scaling) on the log-log plot to highlight this behavior.
Result: Versor demonstrates strictly linear $O(L)$ scaling in both latency and memory usage, as indicated by the parallel slope to the gradient-1 reference line in Figure 3. While Euclidean Transformers experience quadratic memory growth and context-window collapse beyond $T=1000$, Versor maintains structural stability across 10,000 timesteps. With the integration of the compiled C++ Core, the constant-factor overhead of geometric processing is substantially reduced, ensuring that Versor’s $O(L)$ scaling remains efficient even for extremely long-horizon physical trajectories where standard models fail.
@@ 6.2 Generalization Capabilities
Topological Connectivity (“Broken Snake”): We introduce the “Broken Snake” task, adapted from the Pathfinder challenge [17], to test resolution-independent topological reasoning. Models are presented with a grid image containing a snake-like path of activated pixels and must classify whether the path is continuous or broken by a single-pixel gap. Unlike standard CNNs which overfit to feature textures, Versor achieves substantially superior generalization (see Table 2), attaining 99.3% MCC (vs. 50.4% for ViT), because it learns the algebraic law of connectivity (null-displacement vectors) rather than memorizing pixel coordinates.
@@
Broken Snake (Topology) MCC ($\uparrow$) 0.504 (Random) 0.993
Variable System Size ($N=7$) MSE ($\downarrow$) 12.4 (Fail) 3.14
@@
OOD Mass ($10\times$) $\Delta$ Error +3097.2% –19.9%
Variable System Size: Trained on $N=5$, Versor generalizes zero-shot to $N=3,7$ with stable error, while Transformers and HNNs fail due to fixed input dimensions. Hidden Velocity: Without velocity inputs, Versor infers momentum (0.0033 MSE) via recursive state history, whereas the frame-based GATr fails (0.3253 MSE). OOD Mass: On $10\times$ heavier masses, Versor’s error actually improves (-19.9% relative change), while Transformers collapse catastrophically (+3097.2% error increase).
@@ 6.3 Scaling Laws and Capacity
Demonstrating that a novel architecture’s advantages persist across scales is paramount to assessing it potential impact. Here, Versor is evaluated on the Chaotic N-Body Dynamics task across parameter counts ranging from 52k to 901k. The results indicate a consistent "scaling law" behavior where MSE on physical dynamics declines monotonically with the number of parallel multivector channels. Specifically, an error reduction from 25.02 MSE (52k params) to 1.81 MSE (901k params) is observed, following a power-law scaling exponent of $\alpha\approx 0.85$. Unlike standard models where increasing depth often leads to gradient instability, Versor’s manifold-normalized transitions ensure that larger models converge with significantly higher sample efficiency than Euclidean baselines.

### v2

@@ Abstract
A novel sequence architecture is introduced, Versor, which uses Conformal Geometric Algebra (CGA) in place of traditional linear operations to achieve structural generalization and significant performance improvements on a variety of tasks, while offering improved interpretability and efficiency. By embedding states in the $\mathcal{C}l_{4,1}$ manifold and evolving them via geometric transformations (rotors), Versor natively represents $SE(3)$-equivariant relationships without requiring explicit structural encoding.
[…]
@@ 1 Introduction
The remarkable success of the Transformer architecture [31] has cemented the “Sequence of Vectors” paradigm as the dominant framework in artificial intelligence. Whether the modality is text, images (ViT [10]), or audio, data is tokenized and projected into a flat, high-dimensional Euclidean space ($\mathbb{R}^{d_{model}}$), where relationships between features are modeled using the dot product $\mathbf{q}^{T}\mathbf{k}$—a scalar measure of similarity.
However, the physical world is not merely a collection of features but a realization of physical laws acting on a structured manifold. Standard neural networks treat data as points in a flat Euclidean space $\mathbb{R}^{d}$, relying on the dot product as the primary relational primitive. This approach is geometrically naive: it ignores the fundamental symmetries of the physical universe (rotation, translation, and scaling). To respect these symmetries, a standard Transformer must expend vast computational resources to “learn” invariants from millions of augmented examples—a process that approximates what could be analytically enforced by a simple algebraic group action [2]. This fundamental mismatch between the geometry of the world and the architecture of the model is termed the “Euclidean Bottleneck.”
This work argues that the next leap in AI structural intelligence requires embedding these symmetries directly into the substrate of the network. Versor is presented as an architecture built not on linear algebra but on Conformal Geometric Algebra (CGA). Specifically, the architecture operates within $\mathcal{C}l_{4,1}$, a 5-dimensional framework that linearizes the conformal group of 3D Euclidean space.
@@ 1.1 Key Contributions
    Scale Generalization: Demonstration that geometric priors enable scale-invariant reasoning—Versor achieves 0.993 MCC on topological connectivity tasks (vs. 0.070 for Vision Transformers) and maintains structural performance across varying sequence densities.
@@
[…]
@@ 2 Background: Transformers, CGA, and the Manifold Hypothesis
Transformers and the Euclidean Bottleneck. Artificial intelligence is dominated by the Transformer [31]. While effective, its reliance on flat Euclidean space ($\mathbb{R}^{d}$) forces it to learn symmetries ($SE(3)$) from data rather than enforcing them algebraically [2]. This limitation is termed the “Euclidean Bottleneck.”
Conformal Geometric Algebra ($\mathcal{C}l_{4,1}$). The architecture is constructed on $\mathcal{C}l_{4,1}$, a 32-dimensional algebra generated by $\{e_{1},e_{2},e_{3},e_{+},e_{-}\}$. This framework isometrically lifts 3D points $\mathbf{x}$ to null vectors $X$ in 5D space:
@@
This ensures that distance calculations are linearized (see Appendix C.9). Crucially, transformations are represented uniformly as rotors $R$, which act on state vectors $\Psi$ via the sandwich product $\Psi^{\prime}=R\Psi\widetilde{R}$. This structure enforces the Manifold Hypothesis: by constraining latent states to the Spin group, $\text{Spin}(4,1)\subset\mathcal{C}l^{+}_{4,1}$, valid physical transformations (isometries) are explicitly guaranteed, thereby preventing unphysical shearing.
For a detailed introductory treatment of Clifford algebras and their applications, the reader is referred to [8]; see Appendix A for explanations of the basis relations and multivector structure essential to this work.
@@ 3 Related Work
Geometric Deep Learning attempts to encode symmetry priors into neural networks [2], with recent focus on $SO(3)$-equivariant CNNs [6] and the Geometric Algebra Transformer (GATr) [1]. Unlike static-processing baselines such as GATr and CGENN [24], which remain “frame-centric” and require $O(L^{2})$ attention, Versor focuses on path-based sequence modeling. By evolving a latent spinor via the Recursive Rotor Accumulator (RRA), Versor captures temporal coherence as a continuous trajectory on a Lie manifold, providing an inductive bias for dynamical systems that standard frame-based models lack.
This approach aligns with sub-quadratic architectures like Mamba [12] and RWKV [22], but provides a purely geometric interpretation of the hidden state. While concurrent methods like CARE [28] utilize Clifford rotors for positional encodings in Euclidean Transformers, Versor is a complete sequence architecture defined strictly within $\mathcal{C}l_{4,1}$, achieving $O(L)$ linear temporal scaling in its recurrent backbone (though the full network remains $O(L^{2})$ overall due to pairwise attention). Furthermore, by embedding dynamics directly into the manifold, Versor avoids the manually-enforced governing equations required by PINNs [23] and HNNs [11], while achieving better long-horizon stability than GNS [25].
@@ 4 The Versor Architecture
Versor is a sequence architecture that replaces the vector-space assumptions of Transformers with the graded manifold structure of $\mathcal{C}l_{4,1}$. The model processes a stream of multivectors via two core mechanisms: Geometric Product Attention (GPA) for relational reasoning and the Recursive Rotor Accumulator (RRA) for temporal integration. [^1] While the physical dynamics focused on in this work employ $Cl_{4,1}$, the underlying software layout (‘gacore‘) and associated Torch modules naturally support code-generation for Clifford algebras of any parameterised dimension and signature.
@@ 4.2 Component 2: Recursive Rotor Accumulator (RRA)
Manifold Normalization. Crucially, the constraint $\Psi\widetilde{\Psi}=1$ is enforced at every step. This projects numerical drift back onto the manifold, acting as a geometric regularizer that prevents the “exploding state” problem of standard RNNs (see Appendix B for stability proofs).
Hamiltonian Extension. For strict energy conservation, a Hamiltonian-Versor Hybrid is optionally deployed in which the network parameterizes an energy surface $H(q,p)$ rather than predicting steps directly (details in Appendix F.1).
@@ 5 Hardware Acceleration
To address the computational cost of the geometric product ($32^{2}=1024$ operations), two primary execution engines were implemented:
  1. 1.
@@
  2. 2.
    Matrix Isomorphism Acceleration: For the $\mathcal{C}l_{4,1}$ signature, the algebraic isomorphism $\mathcal{C}l_{4,1}\cong\text{Mat}(4,\mathbb{C})$ is leveraged (see Appendix C.7). By mapping multivectors onto this representation space, geometric products are reduced to optimized BLAS GEMM operations, further reducing latency by up to 65% compared to our optimized bitmasked implementation, and by over 95% compared to naive sparse implementations.
@@ 6 Experiments
The Versor architecture is evaluated against state-of-the-art baselines across a variety of challenging tasks: chaotic N-body dynamics, topological reasoning, and multimodal learning (NLP, Vision, Graph). Across all benchmarks, Versor consistently outperforms comparative architectures in accuracy, generalization, and parameter efficiency, validating the effectiveness of geometric inductive biases for sequence modeling. The configurations for each task are summarized in Table 1.
Table 1: Experiment configuration mapping across benchmarks.
Task Configuration
Chaotic N-Body Dynamics RRA (0.007M) / GPA (1.1M)
Long-Horizon N-Body Scaling Pure RRA
Hidden Velocity Inference Pure RRA
Topological Reasoning (Snake) Pure GPA
Molecular Dynamics (MD17) Hybrid (RRA + GPA)
CIFAR-10 Classification Pure GPA
WikiText Language Modeling Pure GPA
@@ 6.1.1 Qualitative Geometric Intelligence
The Geometric Product Attention (GPA) is decomposed into scalar and bivector components (Figure 2). Scalar maps recover distance-based interaction laws, while bivector intensity highlights torque-maximizing interactions, validating that Versor attends to the full geometric configuration of the system.
@@ 6.1.2 Long-Horizon Energy Stability
Models are rolled out for $T=50$ steps. The deviation of total system energy $H=T+V$ (kinetic plus potential) is measured as percentage drift. Conservation of energy is a fundamental physical constraint; a model that has truly learned the underlying dynamics should approximately conserve energy. Parity tests consistently demonstrate that while large-scale models (e.g. 2.2M GATr) can achieve near-perfect conservation, their performance degrades significantly when matched to Versor’s minimal parameter count (6.6K), highlighting the strength of the proposed geometric inductive bias.
Table 2: Performance on Chaotic 5-Body Dynamics (Comparison with State-of-the-Art). Mean $\pm$ std over 5 random seeds. Versor achieves superior accuracy compared to Transformers and general-purpose Graph Networks while maintaining extreme parameter efficiency. Note that HNN achieves superior energy conservation by utilizing hand-coded Hamiltonian energy priors, which Versor matches only in its hybrid “Ham-Versor” configuration without such explicit prior knowledge in the base case.
@@
Transformer ($d=128$) 1.320M 1.10 $6.609\pm 6.415$ $381.1\pm 41.1$
Transformer (Large) 37.8M 12.32 $3.120\pm 1.250$ $210.5\pm 35.2$
@@
GNS [25]* 0.026M 0.29 $5.881\pm 6.408$ $366.7\pm 85.7$
HNN [11] 0.021M 0.11 $4.826\pm 6.378$ $10.7\pm 1.1$
EGNN [26] 0.030M 0.30 $6.695\pm 5.936$ $723.9\pm 351.2$
[…]
* GNS standardized with LayerNorm for stability.
$\ddagger$ Baselines downscaled to match Versor’s parameter scale.
* Latency measured on CPU as per-step inference. $\dagger$ Denotes Bit-Masked (Appendix C.1) vs. Matrix Isomorphism (Appendix C.7) engines. $\ddagger$ Model was downscaled to match Versor’s parameter scale for inductive bias assessment. For fair architectural comparison, GATr was accelerated using the optimized multivector kernel. The 4-channel variant results are detailed in Appendix F.1.4.
[…]
Figure 3: Computational Scaling: Latency (ms) vs. Sequence Length $L$. Versor maintains strictly linear $O(L)$ growth, whereas standard Transformers diverge quadratically, reaching memory exhaustion (OOM) at $L=1024$. The dotted reference line denotes a gradient of 1 (linear scaling) on the log-log plot.
Result. Versor demonstrates strictly linear $O(L)$ scaling in latency, as indicated by the parallel slope to the gradient-1 reference line in Figure 3. While Euclidean Transformers experience quadratic memory growth and context-window collapse beyond $T=1000$, Versor maintains structural stability across 10,000 timesteps. With the compiled C++ Core and Matrix Isomorphism Acceleration, the constant-factor overhead of geometric processing is substantially reduced, ensuring that Versor’s $O(L)$ scaling remains efficient even for extremely long-horizon physical trajectories where standard models fail. Per-step inference latency is reduced to 1.05 ms, outperforming standard $d=128$ Transformers on comparable hardware.
@@ 6.2 Generalization Capabilities
Topological Connectivity (“Broken Snake”). The “Broken Snake” task, adapted from the Pathfinder challenge [18], is introduced to test resolution-independent topological reasoning. Models are presented with a grid image containing a snake-like path of activated pixels and must classify whether the path is continuous or broken by a single-pixel gap. Unlike standard CNNs, which overfit to feature textures, Versor achieves substantially superior generalization (see Table 3), attaining 0.993 MCC (vs. 0.070 for ViT), because it learns the algebraic law of connectivity (null-displacement vectors) rather than memorizing pixel coordinates.
@@
Broken Snake (Topology) MCC ($\uparrow$) 0.070 0.993
[…]
@@
OOD Mass ($10\times m$) $\Delta$ Error +1933.7% –63.9%
[…]
Variable System Size. Trained on $N=5$, Versor generalizes zero-shot to $N=3,7$ with stable error, while Transformers and HNNs fail due to fixed input dimensions.
Hidden Velocity. Without velocity inputs, Versor infers momentum (0.003 MSE) via recursive state history, whereas the frame-based GATr fails (0.3253 MSE).
OOD Mass. On $10\times$ heavier masses, Versor’s error actually improves (-63.9% relative change), while Transformers collapse catastrophically (+1933.7% error increase) [^3] The OOD stress test used a separate controlled evaluation protocol (shorter training horizon of 30 steps) to isolate distribution shift effects. In this setup, the Transformer baseline MSE was 16.16, rising to 328.7 on heavy masses. This differs from the fully-converged reference MSE of 6.61 reported in Table 2.. This counter-intuitive improvement arises because heavier masses increase the system’s inertia, rendering trajectories more momentum-dominated and less chaotic. Versor’s rotors naturally encode momentum conservation, allowing it to exploit this increased predictability. In contrast, Transformers trained on coordinate distributions fail to generalize to the stronger gravitational forces ($F\propto m^{2}$) that drive particles far outside the training distribution.
@@ 6.3 Scaling Laws and Capacity
Demonstrating that a novel architecture’s advantages persist across scales is paramount to assessing its potential impact. Here, Versor is evaluated on the Chaotic N-Body Dynamics task across parameter counts ranging from 52k to 901k. The results indicate consistent “scaling law” behavior, where MSE on physical dynamics declines monotonically with the number of parallel multivector channels. Specifically, an error reduction from $\approx$5.2 MSE (52k params) to 1.82 MSE (901k params) is observed, following a power-law scaling exponent of $\alpha\approx 0.85$. Unlike standard models where increasing depth often leads to gradient instability, Versor’s manifold-normalized transitions ensure that larger models converge with significantly higher sample efficiency than Euclidean baselines.